OpenStudy (rock_mit182):

\int\limits \frac{ (1-x^2)^\frac{ 3 }{ 2 } }{ x^6 }dx I should get - (1-x^2)^(5/2) over 5x^5

OpenStudy (rock_mit182):

\[\int\limits \frac{ (1-x^2)^\frac{ 3 }{ 2 } }{ x^6 }\]

OpenStudy (rock_mit182):

@campbell_st @mathmate

OpenStudy (rock_mit182):

please give me a hand

OpenStudy (anonymous):

Quick tip: wrap your code in `\[ \]` to have it display properly. I would try a trig sub, setting \(x=\sin u\), then \(dx=\cos u\,du\). \[\int\frac{(1-x^2)^{3/2}}{x^6}\,dx=\int\frac{(1-\sin^2u)^{3/2}\cos u}{\sin^6u}\,du=\int\frac{\cos^4u}{\sin^6u}\,du=\int\cot^4u\csc^2u\,du\] Substituting \(t=\cot u\) would be the final step, which gives \(-dt=\csc^2u\,du\).

OpenStudy (anonymous):

Cut-off part is just \(du\).

OpenStudy (rock_mit182):

@SithsAndGiggles according to wolfram the answer is \[\frac{ - (1-x^2)^\frac{ 5 }{ 2 } }{ 5x^5}\]

OpenStudy (anonymous):

What's the problem? Implementing the \(t\) sub gives \[-\int t^4\,dt=-\frac{1}{5}t^5+C=-\frac{1}{5}\cot^5u+C=-\frac{1}{5}\cot^5(\arcsin x)+C\] Are you having trouble showing that this is equal to the answer?